The supplement of an angle measures 40 degres more than twice its complement. Find the measure of the angle
Answer by robertb(5830)

let x be the angle. Then
180 - x = 2(90-x)+40,
180-x = 180-2x+40,
-x = -2x+40,
2x-x = 40, or
x = 40.